The difference equation generated by y=a/x+b


1
Expert's answer
2021-04-14T03:14:15-0400

given: y = "\frac{a}{x}" + b

differentiating y with respect to x we get,

y' = -"\frac{a}{x^2}"


again differentiating y' with respect to x we get,


y'' = "\frac{2a}{x^3}"


"\implies" a = "\frac{y'' x^3}{2}"

putting this value of a in y' we get,

y' = -"\frac{1}{x^2}" "\cdot" "\frac{y'' x^3}{2}"


"\implies" y' = -"\frac{y'' x}{2}"


"\implies" 2y' + y''"\cdot" x = 0

thus the above equation is the differential equation of given y.